Q1 (Process Analysis) - An MSU startup is contemplating a new process to refurbish and recycle rundown equipments. The five steps to the process are:
Step Description Time Required per equipment
1 Register for the process 10.0 minutes
2 Strip off bad parts 15.0 minutes
3 Scrub and clean parts 8.0 minutes
4 Insert new parts 10.0 minutes
5 Polish and deliver 12.0 minutes
One employee will be assigned to each step. Employees will work a 40-hour week (no breaks) and rotate jobs each week. City of Lansing requires 200 refurbished equipments per week but they are unsure that the MSU startup can deliver.
a. What is the processing rate for each step, in Equipments/Hour?
Step Processing Rate (equipments/hour)
1
2
3
4
5
b. What is the bottleneck?
c. Assuming the process runs as designed, what is the maximum weekly output?
d. List three alternative solutions that can achieve throughput of 200 equipments per week and how might they be accomplished? Which of the 3 solutions do you think is the best (and why)?
Answer #1

(a)

Step Description Processing time per unit in minutes
(P)
Number of resources employed
(N)
Processing rate (i.e. capacity) per hr
= 60*N/P
1 Register for the process 10 1 6.0
2 Strip off bad parts 15 1 4.0
3 Scrub and clean parts 8 1 7.5
4 Insert new parts 10 1 6.0
5 Polish and deliver 12 1 5.0

(b)

The bottleneck is the Step-2 i.e. Strip off bad parts for it has the least capacity of 4 equipment per hour

(c)

Maximum output = Capacity of the bottleneck = 4 units per hour = 4*40 units per week = 160 units per week

(d)

The alternatives are as follows:

  • Employ a second operator in step-2. Then the bottleneck will shift 2 step-5 and the capacity will become 5 units per hr. or 200 units per week
  • Give the step-2 operator some incentives so that the processing time improves from 15 minutes to 12 minutes. This will make the capacity 5 units per hr. or 200 units per week
  • Combine step-2 and step-3 as one step and employ 2 operators in that station. Now 'N' will be 2 and 'P' will be 15+8 = 23 minutes. The new station will have capacity = 2*60/23 = 5.21 per hour. Then also the bottleneck will shift 2 step-5 and the capacity will become 5 units per hr. or 200 units per week.

The third solution seems to be the best since both the previous two will require additional costs on a recurring basis. The third solution will only require costs in terms of training the operators which they will already have because of the rotation policy as stated in the question. So, no additional costs.